OpenStudy (anonymous):

Find the x-intercept of 3x + 2y = 12.

OpenStudy (texaschic101):

the x intercept can be found by subbing in 0 for y and solving for x 3x + 2(0) = 12 3x + 0 = 12 3x = 12 x = 12/3 x = 4 so your x intercept is (4,0) any questions ?

OpenStudy (texaschic101):

and the y intercept can be found by subbing in 0 for x and solving for y